# Lift-Attend-Splat: Bird’s-eye-view camera-lidar fusion using transformers

### v1

@@ Abstract
Combining complementary sensor modalities is crucial to providing robust perception for safety-critical robotics applications such as autonomous driving (AD). Recent state-of-the-art camera-lidar fusion methods for AD rely on monocular depth estimation which is a notoriously difficult task compared to using depth information from the lidar directly. Here, we find that this approach does not leverage depth as expected and show that naively improving depth estimation does not lead to improvements in object detection performance and that, strikingly, removing depth estimation altogether does not degrade object detection performance. This suggests that relying on monocular depth could be an unnecessary architectural bottleneck during camera-lidar fusion. In this work, we introduce a novel fusion method that bypasses monocular depth estimation altogether and instead selects and fuses camera and lidar features in a bird’s-eye-view grid using a simple attention mechanism. We show that our model can modulate its use of camera features based on the availability of lidar features and that it yields better 3D object detection on the nuScenes dataset than baselines relying on monocular depth estimation.
@@ 1 Introduction
Integrating information from different modalities efficiently and effectively is especially important in safety-critical applications such as autonomous driving, where different sensor modalities are complementary and combining them adequately is crucial to guarantee safety. For example, cameras capture rich semantic information of objects up to far away distances, while lidars provide extremely accurate depth information but are sparse at large distances. For this reason, many modern self-driving platforms have a large number of different sensors which must be combined together in order to provide accurate and reliable perception of the surrounding scene and allow safe deployment of these vehicles in the real world.
@@
In the context of autonomous driving, many of the recent state-of-the-art methods for camera-lidar fusion [33, 28, 14] are based on the Lift-Splat (LS) paradigm [38] [^1] The “shoot” component of “Lift, Splat, Shoot” [38] relates to trajectory prediction and is not considered here.. In this approach, the camera features are projected in bird’s-eye-view (BEV) — or top-down space — using monocular depth before being fused with the lidar features. As a result, the location of the camera features in BEV is highly dependent on the quality of the monocular depth prediction and it has been argued that its accuracy is critical [28, 14]. In this work we reconsider these claims and show that the monocular depth prediction inside these models is of poor quality and cannot account for their success. In particular, we present results showing that methods based on Lift-Splat perform equally well when the monocular depth prediction is replaced by direct depth estimation from the lidar point cloud or removed completely. This leads us to suggest that relying on monocular depth when fusing camera and lidar features is an unnecessary architectural bottleneck and that Lift-Splat could be replaced by a more effective projection mechanism.
We introduce a novel approach for camera-lidar fusion called “Lift-Attend-Splat” that bypasses monocular depth estimation altogether and instead selects and fuses camera and lidar features in BEV using a simple transformer. We present evidence that our method shows better camera utilisation compared to the methods based on moncular depth estimation and that it improves object detection performance. Our contributions are as follows:
@@ 2 Related work
For 3D object detection, most benchmarks are dominated by methods using lidar point clouds due to their highly accurate range measurement allowing for better placement of objects in 3D compared to methods using cameras or radars only. Deep learning methods for classification on point clouds were pioneered in the seminal works of [40, 41] and early works have been applying similar ideas to 3D object detection [42, 45]. A more recent family of methods is based on direct voxelisation of the 3D space [65, 59] or compression of the lidar representation along the z-direction into “pillars” [22, 60]. These approaches have been very successful and are the basis of many follow-up works [62, 15, 19].
@@
Perception quality can be improved by leveraging jointly cameras and lidars when available. Recent fusion methods can be broadly classified in three categories: point decoration methods, methods that leverage task-specific object queries and architectures, and projection based methods. Point decoration methods augment the lidar point cloud using semantic segmentation data [49, 57], camera features [51], or even create new 3D points using object detections in the image plane [63]. Such methods are relatively easy to implement but suffer from the fact that they require lidar points to fuse camera features. TransFusion [1] is a recent example of a method that leverages task-specific object queries generated using the lidar point cloud. Final detections are made directly without explicit projection of camera features into BEV space. Fusion can also be performed earlier in the model, for example at the level of the 3D voxels [6, 5] or lidar features [23], or by sharing information between the camera and lidar backbones [26, 17, 39]. Finally, projection-based methods project camera features into 3D before fusing them with the lidar (see below).
@@
Of special interest to us are camera-lidar fusion methods based on projecting camera features into 3D. Recent state-of-the-art methods [33, 28, 14] which rank among the top entries in the nuScenes leaderboard [2] leverage the ideas presented in [38] and project camera features in 3D using monocular depth estimation. It was shown in [14] that the performance of [33, 28] can be boosted significantly by including features derived from the ground-truth lidar depth map into the camera stream prior to monocular depth estimation. An alternative approach is to project camera features directly into BEV space using the known correspondence between lidar points and camera features [8, 55, 23]. However, the sparsity of the lidar point cloud can limit which camera features are projected, as described in [33]. Finally, learning to project camera features in BEV without explicit depth can be achieved when lidar is absent using a transformer, as shown in [44, 25]. Here, we extend this line of work to the case of camera-lidar fusion and leverage cross-attention to generate a dense BEV grid of camera features to be fused with the lidar.
CameraRefer to caption LidarRefer to caption Refer to caption BEVFusion [33] Refer to caption BEVFusion [33] w/ $\lambda=1$
@@
No depth 68.5
Figure 1: Impact of the quality of the monocular depth prediction on the object detection performance of BEVFusion [33] on the nuScenes validation set. We compare BEVFusion and four different variants: adding depth supervision using Eq. 2 with various weights $\lambda$, using lidar depth maps instead of monocular depth estimation (lidar), using a pretrained and frozen depth classifier (pretrained), and finally removing depth estimation altogether by projecting camera features at all depths uniformly using Eq. 3 (no depth). In our experiments, more accurate depth does not translate to better detection performance and the original model is on-par with using the lidar points directly as a source of depth. Equivalent detection performance was achieved using the no depth model, clearly indicating that accurate monocular depth is not necessary for BEVFusion [33] to achieve its performance, see main text and Sec. A.4 for details.
@@ 3 Monocular depth prediction in Lift-Splat
Recent camera-lidar fusion methods based on the Lift-Splat paradigm [33, 28] learn a unified representation in the form of a BEV grid by projecting camera features in BEV space using monocular depth estimation as
@@
where ${F^{\prime}}^{\text{cam}}\in\mathbb{R}^{C_{c}^{\prime}\times H\times W}$ is a context vector obtained from the camera features $F^{\text{cam}}\in\mathbb{R}^{C_{c}\times H\times W}$, $D\in\mathbb{R}^{N_{D}\times H\times W}$ is a normalised distribution over predetermined depth bins and Splat denotes the operation of projecting each point downwards into the $z=0$ plane, see [33, 28, 38] for details. The resulting feature map is then merged with the lidar features using concatenation [33] or gated attention [28]. In this paradigm, monocular depth prediction is formulated as a classification problem and learned indirectly from the downstream task without explicit depth supervision.
@@
We analyse the quality of the depth predicted by BEVFusion [33] by comparing it to lidar depth maps, both qualitatively and quantitatively using the absolute relative (Abs. Rel.) and root mean squared errors (RMSE) [9, 24]. As shown in the example displayed on Fig. 1, the depth prediction does not accurately reflect the structure of the scene and is markedly different from the lidar depth map which suggests that monocular depth is not leveraged as expected in [33]. More details can be found in Sec. A.2.
@@
We next investigate whether improving the the depth prediction quality can boost object detection performance. To do so, we retrain the model from [33] with the following loss:
@@
where $L_{\text{sup}}$ is the original 3D object detection loss and $L_{\text{depth}}$ is a simple cross-entropy loss for the depth estimation that uses the lidar depth as a target, see Sec. A.3 for more details. By changing the hyper-parameter $\lambda$, we can control the quality of the depth prediction and explore how it impacts detection performance. In Fig. 1, we see that while depth supervision indeed leads to much more accurate depth maps both visually and quantitatively, detection performance — measured using mean average precision (mAP) — degrades from the baseline as the weight of the depth supervision is increased. This suggests that the method is unable to take advantage of more accurate depth prediction. Since training on the multi-task loss Eq. 8 is likely to degrade object detection performance at high values of $\lambda$, we also experiment with two more variants: (i) pretraining the depth supervision module separately and (ii) using the lidar point cloud directly to bypass the depth supervision module altogether. Pretraining leads to more accurate depth prediction but degrades detection performance relative to the baseline, while using the lidar directly does not change the detection performance compared to the baseline, even though all depth metrics are close to zero [^2] They are not exactly zero because of the depth quantisation introduced by the one-hot encoding of the lidar depth, see Sec. A.1..
@@ 4 Camera-lidar fusion without monocular depth estimation
[width=] settings.render=6; settings.prc=false; import three; import graph3; import grid3; currentprojection=perspective(6, 3, 2);
//Draw Axes pen thickblack = black+0.75;
//Set parameters of start corner of polar volume element real r = 1; real f=pi/4; //phi
real df=pi/16; //dphi real dr=0.1;
pen thinblack = black+0.25; pen thingrey = grey+0.25;
real gridZ = -0.5;
real minGridX = -0.2; real maxGridX = 0.8; real minGridY = 0.1; real maxGridY = 1.1; real gridDimX = 10; real gridDimY = 10; // grid for(int i=0; i¡=gridDimX; ++i) real x = minGridX + ((i / gridDimX) * (maxGridX - minGridX)); draw((x,minGridY,gridZ)–(x,maxGridY,gridZ), thingrey); for(int i=0; i¡=gridDimY; ++i) real y = minGridY + ((i / gridDimY)* (maxGridY - minGridY)); draw((minGridX,y,gridZ)–(maxGridX,y,gridZ), thingrey);
// camera grid real minGridX = -.3; real maxGridX = .3; real gridY = -0.3; real minGridZ = -0.2; real maxGridZ = .2; real gridDimX = 8; real gridDimZ = 5;
// grid real deltaX = (maxGridX - minGridX) / gridDimX; for(int i=0; i¡=gridDimX; ++i) real x = minGridX + i * deltaX; draw((x,gridY,minGridZ)–(x,gridY,maxGridZ), thinblack); real deltaZ = (maxGridZ - minGridZ) / gridDimZ; for(int i=0; i¡=gridDimZ; ++i) real z = minGridZ + i * deltaZ; draw((minGridX,gridY,z)–(maxGridX,gridY,z), thinblack); path3 col = (minGridX + 1*deltaX,gridY, minGridZ)–(minGridX + 2*deltaX,gridY,minGridZ)–(minGridX + 2*deltaX,gridY,maxGridZ)–(minGridX + 1*deltaX,gridY,maxGridZ)–cycle; draw(col, blue+0.75);
real phiMinY = 0.3; real phiMaxY = 1.0;
//phi arcs for (int i=0; i ¡ 9; ++i) real startPhiX = -phiMinY + (2 * phiMinY * i / 8); real endPhiX = -phiMaxY + (2 * phiMaxY * i / 8); draw((startPhiX, phiMinY, 0)–(endPhiX, phiMaxY, 0), thinblack);
// horizon line draw((minGridX,gridY,0)–(maxGridX,gridY,0), dashed+thinblack);
// horizon edges draw((gridY,gridY,0)–(phiMinY,phiMinY,0), dashed+thinblack); draw((-gridY,gridY,0)–(-phiMinY,phiMinY,0), dashed+thinblack);
// The focal point! dot((0,0,0),black);
for (int i=0; i ¡ 8; ++i) real y = phiMinY + (i * (phiMaxY - phiMinY) / 7); draw((-y, y, 0)–(y, y, 0), thinblack);
real closeMinX = -phiMinY + (2 * phiMinY * 6 / 8); real closeMaxX = -phiMinY + (2 * phiMinY * 7 / 8); real farMinX = -phiMaxY + (2 * phiMaxY * 6 / 8); real farMaxX = -phiMaxY + (2 * phiMaxY * 7 / 8);
draw((closeMinX,phiMinY,0)–(closeMaxX,phiMinY,0)–(farMaxX,phiMaxY,0)–(farMinX,phiMaxY,0)–cycle, red+0.75);
//dashed connectors real dashedFrac = 6.5; real dashedDY = 0.1; int numDashes = 7; for(int i = 0; i ¡ numDashes; ++i) real y = phiMinY + ((phiMaxY - phiMinY) * (i + 0.5) / numDashes); real x = -y + (2 * y) * (dashedFrac / gridDimX); draw((x, y, gridZ)–(x, y, 0), dotted, arrow=Arrow3(TeXHead2, Relative(.6))); draw((x, y, 0)–(x, y, gridZ), dotted, arrow=Arrow3(TeXHead2, Relative(.8)));
@@
Figure 2: Lift-Attend-Splat camera-lidar fusion architecture. (left) Overall architecture: features from the camera and lidar backbones are fused together and merged before being passed to a detection head. (inset) Geometry of our 3D projection: the “Lift” step embeds the lidar BEV features into the projected horizon by lifting the lidar features along the z-direction using bilinear sampling. The “Splat” step corresponds to the inverse transformation in that it projects features from the projected horizon back onto the BEV grid using bilinear sampling, again along the z-direction. (right) Details of the projection module: the “Attend” step in our method let the lifted lidar features $\tilde{B}^{\text{lid}}_{i}$ attend to the camera features $F^{\text{cam}}_{i}$ in the corresponding column using a simple encoder-decoder transformer architecture to produce fused features $D(\tilde{B}_{i}^{\text{lid}},E(F^{\text{cam}}_{i}))$ in frustum space.
In this section, we present a camera-lidar fusion method that bypasses monocular depth estimation altogether and instead fuses camera and lidar features in bird’s-eye-view using a simple transformer [48]. A naive application of the transformer architecture to the problem of camera-lidar fusion is difficult, however, due to the large number of camera and lidar features and the quadratic nature of attention. As shown in [44], it is possible to use the geometry of the problem to drastically restrict the scope of the attention when projecting camera features in BEV, since camera features should only contribute to locations along their corresponding rays. We adapt this idea to the case of camera-lidar fusion and introduce a simple fusion method that uses cross-attention between columns in the camera plane and polar rays in the lidar BEV grid. Instead of predicting monocular depth, the cross-attention learns which camera features are the most salient given context provided by the lidar features along its ray.
With the exception of the projection of the camera features in BEV, our model shares a similar overall architecture to methods based on the Lift-Splat paradigm [33, 28, 14] and is depicted on Fig. 2 left. It consists of the following modules: the camera and lidar backbones which produce features for each modality independently, a projection and fusion module that embeds the camera features into BEV and fuses them with the lidar, and finally a detection head. When considering object detection, the final output of the model is the property of objects in the scene represented as 3D bounding boxes with position, dimension, orientation, velocity and classification information. In what follows we explain in detail the architecture of our projection and fusion modules.
@@
where $\mathbf{C}$ is the $3\times 4$ camera projection matrix (intrinsic and extrinsic), and $h$ is the height of the image. Note that this plane is not in general parallel to the BEV grid, its relative orientation being defined by the camera’s extrinsic parameters. We define a regular grid on the projected horizon that is aligned with the 2D grid of features in the image plane by tracing out rays from the intersection of the horizontal line with the edges of the feature columns in the image plane, and then separating these rays into a set of predetermined depth bins (similarly to [28]). Features on this grid can be represented by a matrix $G\in\mathbb{R}^{N_{D}\times W}$, where each row corresponds to a specific column in the camera feature map $F^{\text{cam}}\in\mathbb{R}^{H\times W\times C}$. The geometry of the projected horizon can be seen in Fig. 2 (left, inset).
@@
We can easily define a correspondence between points on the projected horizon and points on the BEV plane by projecting them along the z-direction in 3D space. As cameras are in general tilted with respect to the ground, this correspondence depends on each camera’s extrinsic parameters. We transfer lidar features from the BEV grid to a camera’s projected horizon through bi-linear sampling of the BEV grid at the locations of the down-projected cell-centers of the projected horizon. We call this process “lifting” and denote it as $\text{Lift}_{i}$ for the projected horizon of camera $i$. Similarly, features can be transferred in the opposite direction, from projected horizon to BEV grid, by bi-linearly sampling the projected horizon at the locations of the projected cell-centers of the BEV grid. We denote this operation as $\text{Splat}_{i}$, similarly to [38, 33, 28].
@@
Our projection module is depicted in Fig. 2 (right) and can be broken down into three simple steps: (i) we first lift the BEV lidar features $B^{\text{lid}}$ onto the projected horizon of camera $i$, producing “lifted” lidar features $\tilde{B}^{\text{lid}}_{i}$, (ii) we then let the “lifted” lidar features attend to the camera features in the corresponding column using a simple transformer encoder-decoder, producing fused features $\tilde{B}^{\text{fus}}_{i}$ on the projected horizon, and finally (iii) we splat these features back onto the BEV grid to produce $B^{\text{fus}}_{i}$. During the attend step, the camera features in each column are encoded by a transformer encoder $E$ and passed as keys and values to a transformer decoder $D$ which uses the frustum lidar features as queries. The result of these three steps can be written as
@@
where $\text{Lift}_{i}$ and $\text{Splat}_{i}$ project the BEV features onto the projected horizon of camera $i$ (and vice versa) as described above. Finally, we apply a simple fusion module where we sum the projected features from different cameras together, concatenate them with the lidar features and apply a convolutional block to obtain the final features in BEV. This simple architecture allows the camera features to be projected from the image plane onto the BEV grid without requiring monocular depth estimation. We use a single set of column-frustum transformer weights which are shared across all column-frustum pairs and cameras. For simplicity, we use here a single transformer encoder and decoder but show that increasing the number of such blocks can be beneficial in Sec. 5.4.
@@
It is worth discussing how our approach differs from predicting monocular depth directly. When using monocular depth, each feature in the camera feature map is projected into BEV at multiple locations weighted by a normalised depth distribution. This normalisation limits each feature to be projected either into a single location or smeared with lower intensity across multiple depths. However, in our approach, the attention between camera and lidar is such that the same camera feature can contribute fully to multiple locations in the BEV grid. This is possible because attention is normalised over keys, which correspond to different heights in the camera feature map, rather than queries, which correspond to different distances along the ray. Furthermore, our model has access to lidar features in BEV when choosing where to project camera features, which gives it greater flexibility.

### v3

@@ Abstract
Combining complementary sensor modalities is crucial to providing robust perception for safety-critical robotics applications such as autonomous driving (AD). Recent state-of-the-art camera-lidar fusion methods for AD rely on monocular depth estimation which is a notoriously difficult task compared to using depth information from the lidar directly. Here, we find that this approach does not leverage depth as expected and show that naively improving depth estimation does not lead to improvements in object detection performance. Strikingly, we also find that removing depth estimation altogether does not degrade object detection performance substantially, suggesting that relying on monocular depth could be an unnecessary architectural bottleneck during camera-lidar fusion. In this work, we introduce a novel fusion method that bypasses monocular depth estimation altogether and instead selects and fuses camera and lidar features in a bird’s-eye-view grid using a simple attention mechanism. We show that our model can modulate its use of camera features based on the availability of lidar features and that it yields better 3D object detection on the nuScenes dataset than baselines relying on monocular depth estimation.
@@ 1 Introduction
Integrating information from different modalities efficiently and effectively is especially important in safety-critical applications such as autonomous driving, where different sensor modalities are complementary and combining them adequately is crucial to guarantee safety. For example, cameras capture rich semantic information of objects up to far away distances, while lidars provide extremely accurate depth information but are sparse at large distances. For this reason, many modern self-driving platforms have a large number of different sensors which must be combined in order to provide accurate and reliable perception of the surrounding scene and allow safe deployment of these vehicles in the real world.
@@
In the context of autonomous driving, many of the recent state-of-the-art methods for camera-lidar fusion [36, 31, 16] are based on the Lift-Splat (LS) paradigm [41] [^1] The “shoot” component of “Lift, Splat, Shoot” [41] relates to trajectory prediction and is not considered here.. In this approach, the camera features are projected in bird’s-eye-view (BEV) — or top-down space — using monocular depth before being fused with the lidar features. As a result, the location of the camera features in BEV is highly dependent on the quality of the monocular depth prediction and it has been argued that its accuracy is critical [31, 16]. In this work, we reconsider these claims and show that the monocular depth prediction inside these models is of poor quality and cannot account for their success. In particular, we present results showing that methods based on Lift-Splat perform equally well when the monocular depth prediction is replaced by direct depth estimation from the lidar point cloud or removed completely. This leads us to suggest that relying on monocular depth when fusing camera and lidar features is an unnecessary architectural bottleneck and that Lift-Splat could be replaced by a more effective projection.
We introduce a novel approach for camera-lidar fusion called “Lift-Attend-Splat” that bypasses monocular depth estimation altogether and instead selects and fuses camera and lidar features in BEV using a simple transformer. We present evidence that our method shows better camera utilisation compared to the methods based on monocular depth estimation and that it improves object detection performance. Our contributions are as follows:
@@ 2 Related work
Most 3D object detection benchmarks are dominated by methods using lidar point clouds due to their highly accurate range measurement allowing for better placement of objects in 3D compared to methods using cameras or radars only. Deep learning methods for classification on point clouds were pioneered in the seminal works of [43, 44] and early works have been applying similar ideas to 3D object detection [45, 48]. A more recent family of methods is based on direct voxelisation of the 3D space [70, 64] or compression of the lidar representation along the z-direction into “pillars” [25, 65]. These approaches have been very successful and are the basis of many follow-up works [67, 17, 22].
@@
Perception quality can be improved by jointly leveraging cameras and lidars when available. Recent fusion methods can be broadly classified into three categories: point decoration methods, methods that leverage task-specific object queries and architectures, and projection based methods. Point decoration methods augment the lidar point cloud using semantic segmentation data [52, 61], camera features [54], or even create new 3D points using object detections in the image plane [68]. Such methods are relatively easy to implement but suffer from the fact that they require lidar points to fuse camera features. TransFusion [1] is a recent example of a method that leverages task-specific object queries generated using the lidar point cloud. Final detections are made directly without explicit projection of camera features into BEV space. Similar methods utilising two-way modality interactions perform even better [66, 62]. Fusion can also be performed earlier in the model, for example at the level of the 3D voxels [7, 6] or lidar features [26], or by sharing information between the camera and lidar backbones [29, 19, 42]. MSMDFusion [21] fuses camera and lidar features at multiple scales using lidar points to estimate the 3D position of camera features, while UniTR [55] pre-assigns depth to each camera feature. FUTR3D [5] fuses features by selecting modality-agnostic 3D reference points. Finally, projection-based methods project camera features into 3D before fusing them with the lidar (see below).
@@
Of most interest to us are camera-lidar fusion methods based on projecting camera features into 3D. Many recent state-of-the-art camera-lidar fusion methods [36, 31, 16] project camera features in 3D using monocular depth estimation even though depth information is available from the lidar. It has also been shown that the projection method is less important than other aspects of training in the camera-only setting [14]. An alternative approach is to project camera features directly into BEV space using the known correspondence between lidar points and camera features [9, 59, 26]. However, the sparsity of the lidar point cloud can limit which camera features are projected, as described in [36]. Finally, learning to project camera features in BEV without explicit depth can be achieved when lidar is absent using a transformer, as shown in [47, 28]. Here, we extend this line of work to the case of camera-lidar fusion and leverage cross-attention to generate a dense BEV grid of fused lidar features.
Camera
Lidar
BEVFusion [36]
BEVFusion [36] w/ $\lambda=1$
@@
Uniform depth 68.5
Figure 1: Impact of the quality of the monocular depth prediction on the object detection performance of BEVFusion [36] on the nuScenes validation set. We compare BEVFusion and four different variants: adding depth supervision using Eq. 2 with various weights $\lambda$, using lidar depth maps instead of monocular depth estimation (lidar), using a pretrained and frozen depth classifier (pretrained), and finally removing depth estimation altogether by projecting camera features at all depths uniformly using Eq. 3 (uniform depth). In our experiments, more accurate depth does not translate to better detection performance and the original model is on-par with using the lidar points directly as a source of depth. Equivalent detection performance was achieved using the uniform depth model, clearly indicating that accurate monocular depth is not necessary for BEVFusion [36] to achieve its performance, see main text and Sec. A.4 for details.
@@ 3 Monocular depth prediction in Lift-Splat
Recent camera-lidar fusion methods based on the Lift-Splat paradigm [36, 31] learn a unified representation in the form of a BEV grid by projecting camera features in BEV space using an estimated depth distribution as
@@
where ${F^{\prime}}^{\text{cam}}\in\mathbb{R}^{C_{c}^{\prime}\times H\times W}$ is a context vector obtained from the camera features $F^{\text{cam}}\in\mathbb{R}^{C_{c}\times H\times W}$, $D\in\mathbb{R}^{N_{D}\times H\times W}$ is a normalised distribution over predetermined depth bins and Splat denotes the operation of projecting each point downwards into the $z=0$ plane, see [36, 31, 41] for details. The resulting feature map is then merged with the lidar features using concatenation [36] or gated attention [31]. In this paradigm, the monocular depth distribution prediction is learned indirectly from the downstream task without explicit depth supervision.
@@
We analyse the quality of the depth distribution predicted by BEVFusion [36] by comparing its mean value to lidar depth maps, both qualitatively and quantitatively using the absolute relative (Abs. Rel.) and root mean squared errors (RMSE) [10, 27]. As shown on Fig. 1, the mean depth prediction does not accurately reflect the structure of the scene and is markedly different from the lidar depth map which suggests that monocular depth is not leveraged as expected in [36]. See Secs. A.2 and A.4 for details.
@@
We next investigate whether improving the depth prediction quality can boost object detection performance. To do so, we retrain the model from [36] with the following loss:
@@
where $L_{\text{sup}}$ is the original 3D object detection loss and $L_{\text{depth}}$ is a simple cross-entropy loss for the depth estimation that uses one-hot encoded lidar depth as a target, see Sec. A.3 for more details. By changing the hyper-parameter $\lambda$, we can control the quality of the depth prediction and explore how it impacts detection performance. In Fig. 1, we see that while depth supervision indeed leads to much more accurate depth maps both visually and quantitatively, detection performance — measured using mean average precision (mAP) — degrades from the baseline as the weight of the depth supervision is increased. This suggests that the method is unable to take advantage of more accurate depth prediction. Since training on the multi-task loss Eq. 7 is likely to degrade object detection performance at high values of $\lambda$, we also experiment with two more variants: (i) pretraining the depth supervision module separately and (ii) using the lidar point cloud directly to bypass the depth supervision module altogether. Pretraining leads to more accurate depth prediction but degrades detection performance relative to the baseline, while using the lidar directly does not change the detection performance compared to the baseline, even though all depth metrics are close to zero [^2] They are not exactly zero because of the depth quantisation introduced by the one-hot encoding of the lidar depth, see Sec. A.1..
@@ 4 Camera-lidar fusion without monocular depth estimation
LiftAttendSplatcFusionProjection Lidar encoder Camera encoder Refer to captionRefer to captionRefer to captionRefer to captionDetection headRefer to captionImage planeProjected horizonBEV grid
@@
Figure 3: Lift-Attend-Splat camera-lidar fusion architecture. (left) Overall architecture: features from the camera and lidar backbones are fused together and merged before being passed to a detection head. (inset) Geometry of our 3D projection: the “Lift” step embeds the lidar BEV features into the projected horizon by lifting the lidar features along the z-direction using bilinear sampling. The “Splat” step corresponds to the inverse transformation in that it projects features from the projected horizon back onto the BEV grid using bilinear sampling, again along the z-direction. (right) Details of the projection module: the “Attend” step in our method lets the lifted lidar features $\tilde{B}^{\text{lid}}_{i}$ attend to the camera features $F^{\text{cam}}_{i}$ in the corresponding column using a simple encoder-decoder transformer architecture to produce fused features $D(\tilde{B}_{i}^{\text{lid}},E(F^{\text{cam}}_{i}))$ in frustum space.
In this section, we present a camera-lidar fusion method that bypasses monocular depth estimation altogether and instead fuses camera and lidar features in bird’s-eye-view using a simple transformer [51]. A naive application of the transformer architecture to the problem of camera-lidar fusion is difficult, however, due to the large number of camera and lidar features and the quadratic nature of attention. As shown in [47], it is possible to use the geometry of the problem to drastically restrict the scope of the attention when projecting camera features in BEV, since camera features should only contribute to locations along their corresponding rays. We adapt this idea to the case of camera-lidar fusion and introduce a simple method that uses cross-attention between columns in the camera plane and polar rays in the lidar BEV grid. Instead of predicting monocular depth, cross-attention learns which camera features are the most salient given context provided by the lidar features along its ray.
Except for the projection of the camera features in BEV, our model shares a similar overall architecture to methods based on the Lift-Splat paradigm [36, 31, 16] and is depicted on Fig. 3 left. It consists of the following modules: the camera and lidar backbones which produce features for each modality independently, a projection and fusion module that embeds the camera features into BEV and fuses them with the lidar, and finally a detection head. When considering object detection, the final output of the model is the property of objects in the scene represented as 3D bounding boxes with position, dimension, orientation, velocity and classification information. In what follows we explain in detail the architecture of our projection and fusion modules.
@@
where $\mathbf{C}$ is the $3\times 4$ camera projection matrix (intrinsic and extrinsic), and $h$ is the height of the image. Note that this plane is not in general parallel to the BEV grid, its relative orientation being defined by the camera’s extrinsic parameters. We define a regular grid on each camera’s projected horizon that is aligned with the 2D grid of features in its image plane by tracing out rays from the intersection of the horizontal line with the edges of the feature columns in the image plane, and then separating these rays into a set of predetermined depth bins (similarly to [31]). Features on this grid can be represented by a matrix $G\in\mathbb{R}^{N_{D}\times W}$, where each row corresponds to a specific column in the camera feature map $F^{\text{cam}}\in\mathbb{R}^{H\times W\times C}$. The geometry of a projected horizon can be seen in Fig. 3 (left, inset). The projected horizon allows for a consistent definition of depth between differently pitched cameras.
@@
We can easily define a correspondence between points on a projected horizon and points on the BEV plane by projecting them along the z-direction in 3D space. As cameras are in general tilted with respect to the ground, this correspondence depends on each camera’s extrinsic parameters. We transfer lidar features from the BEV grid to a camera’s projected horizon through bi-linear sampling of the BEV grid at the locations of the down-projected cell-centers of the projected horizon. We call this process “lifting” and denote it as $\text{Lift}_{i}$ for the projected horizon of camera $i$. Similarly, camera features can be transferred in the opposite direction, from a projected horizon to the BEV grid, by bi-linearly sampling the projected horizon at the locations of the projected cell-centers of the BEV grid. We denote this operation as $\text{Splat}_{i}$, similarly to [41, 36, 31]. Fusion of lidar features with splatted camera features takes place in BEV space, as is common [36, 31].
@@
Our projection module is depicted in Fig. 3 (right) and can be broken down into three simple steps: (i) we first lift the BEV lidar features $B^{\text{lid}}$ onto the projected horizon of camera $i$, producing “lifted” lidar features $\tilde{B}^{\text{lid}}_{i}$, (ii) we then let the “lifted” lidar features attend to the camera features in the corresponding column using a simple transformer encoder-decoder, producing fused features $\tilde{B}^{\text{fus}}_{i}$ on the $i^{\text{th}}$ projected horizon, and finally (iii) we splat these features back onto the BEV grid to produce $B^{\text{fus}}_{i}$. During the attend step, the camera features in each column are encoded by a transformer encoder $E$ and passed as keys and values to a transformer decoder $D$ which uses the frustum lidar features as queries. The result of these three steps can be written as
@@
where $\text{Lift}_{i}$ and $\text{Splat}_{i}$ project the BEV features onto the projected horizon of camera $i$ (and vice versa) as described above. Finally, we apply a simple fusion module where we sum the projected features from different cameras together, concatenate them with the lidar features and apply a convolutional block to obtain the final features in BEV. This simple architecture allows the camera features to be projected from the image plane onto the BEV grid without requiring monocular depth estimation. We share a single set of transformer weights across all column-frustum pairs and cameras. For simplicity, we use here a single transformer encoder and decoder but show in Sec. 5.4 that adding more can be beneficial. All camera features participate in our attention, not just a small number of reference points as is the case in [28]. This learnt set of salient features initialises our object detection queries, rather than the fixed maxpool of [1].
@@
It is worth discussing how our approach differs from predicting monocular depth directly. When using monocular depth, each feature in the camera feature map is projected into BEV at multiple locations weighted by a normalised depth distribution. This normalisation limits each feature to be projected either into a single location or smeared with lower intensity across multiple depths. However, in our approach, the attention between camera and lidar is such that the same camera feature can contribute fully to multiple locations in the BEV grid. This is possible because attention is normalised over keys, which correspond to different heights in the camera feature map, rather than queries, which correspond to different distances along the ray. Furthermore, our model has access to lidar features in BEV when choosing where to project camera features, which gives it greater flexibility. Finally, our projection requires fewer parameters than competing methods: 0.9M for our attention-based module compared to 1.6M in the equivalent component of [36].